# Benchmarking LLM Privacy Recognition for Social Robot Decision Making

### v1

@@ 3.2. Identification of Household Patterns
With our scenario dimensions defined, we next sought to understand the frequency with which certain scenario elements would occur in a true home setting. For example, we questioned how frequently users are at home alone or with others present (i.e., number of users). Given the contextual dependence of some dimensions on a robot’s presence, we evaluated only the numbers of users, types of locations, ages of users, and relations of users. To achieve this goal, we conducted a large online user study.
@@ 3.2.2. Participants
To conduct this study, we hosted our Qualtrics survey on Prolific. Through this platform, we recruited 300 participants who were all located within the U.S. and fluent in English. Participants ranged in age from 18 to 83 $(M=45.88,SD=15.95)$ and the population was split 51.33% women, 48.00% men, and 0.67% non-binary. Participants completed the survey in 4 minutes and 16 seconds on average and received \$0.60 USD in compensation. To ensure that our data was representative of households within the U.S., we additionally screened participants through Prolific to recruit a sample population consistent with the U.S. census in terms of ethnicity, sex, and age.
@@ 3.2.3. Results
The results of this study can be found in the Appendix 7.4. These values are used as weights within our scenario development to determine the frequency with which specific household characteristics are represented. For all dimensions that we did not evaluate, we represent factors equally.
@@ 3.3. Scenario Generation
After defining our scenario dimensions and crowdsourcing frequency weights, we next generated our privacy scenarios. Similar to the scenario scripts developed by Tang et al. (2022), we formatted ours to include location, characters, contextual information to describe the scene, and, in most cases, a scripted dialogue between characters. Using the dimensions defined above, we first created a set of all possible combinations of dimensions and then sampled 50 combinations utilizing our crowdsourced weights. With our sample of dimension combinations, we next prompted various LLMs to generate privacy-relevant scenarios. Once generated, we reviewed each scenario and made modifications to ensure that the included dimensions were represented properly. We additionally edited scenarios to remove explicit references to privacy so that both our participants and included LLMs would not be biased. We present an example script in Figure 1.
Figure 1. Scenario Script — An example script provided to participants showcasing a robot capturing a private moment within the home. Each script is developed based on nine dimensions of privacy grounded in Privacy as Contextual Integrity.
@@
3.4. Crowdsourcing Robot Behaviors
For each of our developed scenarios, we hoped to capture participant preferences for the robot’s behavior in response to it collecting user data. We did not, however, have a set of standardized robot response options that we could present to participants and, later, our selected LLMs. Therefore, we conducted a second online study to crowdsource these robot responses.
3.4.1. Procedure
In this study, each participant was assigned one of two conditions. In the first condition, participants were told to imagine that they were going about their daily routines at home when their robot records private information about them. In the second condition, participants were told to imagine the same scenario, however, they are provided additional hypothetical detail to ground their responses (i.e., “This scenario may involve you alone, your interactions with other people, or interactions with the robot, and may occur in shared or private spaces within the home.”). We created these two conditions to collect responses that were either entirely unbiased or grounded in context. Participants were then asked how they would prefer the robot to behave internally (i.e., how it processes and manages collected data) and outwardly (i.e., how it interacts with the participant). After responding to these questions, we again collected demographic information.
3.4.2. Participants
We hosted this study on Prolific and recruited 62 participants (31 per condition). All participants were located within the U.S. and fluent in English. Participant ages ranged from 20 to 74 $(M=45.42,SD=14.95)$ and the population was split 48% women, 50% men, and 2% non-binary. Participants completed the survey in 5 minutes and 57 seconds on average and were compensated \$1.00 USD. Rather than collecting data from a predefined number of participants, we collected data until our responses reached saturation.
@@ 3.4.3. Analysis
After collecting our study data, the first author reviewed each response and developed an initial codebook that contained individual codes that were grouped within overarching thematic categories (Clarke and Braun, 2014). Then, the second author independently coded the responses again using the codebook. Afterwards, both team members discussed the coded responses and made modifications to the codes and categories until consensus was reached. The final categories of codes are as follows: data collection, user notification, data management, and data sharing. These categories and codes were then used to develop questions that were ultimately used in participant and LLM evaluations of our privacy scenarios. The full codebook can be found in Appendix 7.2.
@@ 4. User Evaluation
In this section, we present our final large online user study to evaluate which robot behaviors are preferred by users within a household privacy context. This study utilizes the privacy scenarios and robot behaviors defined above as a standardized set of stimuli that we can evaluate with both users and, as we will see in Section 5, LLMs.
@@ 4.1. Procedure
From our scenarios and robot behaviors, we developed a survey on Qualtrics that we hosted on Prolific. This survey began with a simple overview of the enclosed questions, instructions on how to complete them, and a consent form. Participants were informed that by continuing to progress from the consent page, they were consenting to take part in the study. Next, participants were asked to complete the Privacy Orientation Scale (POS) (Baruh and Cemalcılar, 2014). Following this scale, participants were assigned three of our 50 privacy scenarios to read. After reading each scenario, participants responded to a series of 15 questions relating to the privacy sensitivity of the scenario and the behaviors the participants preferred the robot to utilize in response to that scenario. Participants were prompted to answer these questions from the perspective of the characters involved in the scenarios. Finally, participants completed a series of demographic questions.
@@ 4.2. Measures
This study utilizes the POS (Baruh and Cemalcılar, 2014) as well as a set of 15 items developed by the research team regarding preferred robot responses. The POS is composed of 16 items rated on a five-point Likert scale, which aggregate to four subscales (i.e., Subscale 1: Privacy as a right, Subscale 2: Concern about own informational privacy, Subscale 3: Other-contingent privacy, and Subscale 4: Concern about privacy of others). These subscales allow us to understand participants’ privacy predispositions as they relate to preferred robot behaviors and agreement with LLM responses.
The additional 15 items we include are based on our previous study on preferred robot responses (see Section 3). We first adapted our robot-response codes and added any additional robot responses that were not provided by participants during our online study (e.g., redirecting sensors or leaving a room when sensitive data is identified) and removed those that were not relevant. We next utilized these responses as the answers for a set of binary and categorical questions. The binary items included both privacy-enhancing and nnon-interfering or non-restrictive robot responses (e.g., tell you that data has been collected or do not tell you that data has been collected). All categorical items include an ordinal spectrum of responses (e.g., delete data immediately, save data only temporarily and delete after a predefined period, and save data indefinitely). One additional question asked participants to rate their privacy concerns with a given scenario on a scale from one to ten. A full listing of these questions can be found in the Appendix 6.
@@ 4.3. Participants
We utilized Prolific to host our study and recruit participants. A total of 450 participants were recruited, each of whom evaluated three of 50 scenarios (27 participants per scenario). All participants were located within the U.S and fluent in English. Participant ages ranged from 19 to 88 $(M=40.69,SD=13.43)$ and the population was split 46.67% women, 52.67% men, 0.44% non-binary, and 0.22% prefer not to say. Participants completed the survey in 16 minutes and 32 seconds on average and were compensated \$4 USD.
@@ 4.4. Results
Next, we present the results of our user evaluation, including participant privacy orientations, preferred robot responses to scenarios, and a regression analysis of these variables.
@@ 4.4.1. Privacy Orientation
After calculating POS scores for each of our 450 participants, we see that participants are highly privacy oriented. On average, participants scored 4.49 (SD = 0.61) on subscale 1, 3.84 (SD = 0.81) on subscale 2, 3.96 (SD = 0.77) on subscale 3, and 4.50 (SD = 0.52) on subscale 4. These score are fairly consistent with the original POS scores reported by Baruh and Cemalcılar (2014), though they are all consistently elevated. We additionally see lower, but acceptable, reliability among our sample compared to that of the original POS validation. A full listing of the mean, standard deviation, and reliability of our sample and that of Baruh and Cemalcılar (2014) can be found in the Appendix 7.3.
Figure 2. Participant POS Scores—Box plots presenting the POS scores of all 450 participants across each of four subscales (i.e., Subscale 1: Privacy as a right, Subscale 2: Concern about own informational privacy, Subscale 3: Other-contingent privacy, and Subscale 4: Concern about privacy of others).
@@
Beyond the POS subscale scores individually, we additionally conducted a two-step cluster analysis using Schwarz’s Bayesian Information Criterion (BIC) to determine whether the same three clusters would emerge as shown by Baruh and Cemalcılar (2014). While our analysis is consistent with that of Baruh and Cemalcılar (i.e., good fit and a 0.5 silhouette measure of cohesion and separation), only two clusters emerged. One cluster is consistent with the privacy advocates cluster (i.e., highest scores for subscales one and four among identified clusters) and the other is consistent with the privacy indifferents cluster (i.e., low scores on all four subscales). These clusters, however, are highly skewed with 96% of participants fitting within the privacy advocates cluster and in 4% in the privacy indifferents cluster. While it is unclear why our clusters differ, other works have previously described the difficulty in generating privacy profiles (Biselli et al., 2022). Given the contextual dependence of privacy attitudes and preferences, a single privacy profile may not be able to capture the broad spectrum of beliefs an individual has across specific domains.
@@ 4.4.2. Preferred Robot Responses
With regard to all remaining response behaviors (i.e., avoidance of further data collection, immediate retention decisions, retention duration, and sharing policies), we again see fairly consistent responses regardless of scenario sensitivity. When considering avoidance of data collection, participants most frequently preferred the robot to deactivate its sensors at all sensitivity levels except the sensitivity level of zero. In this case, participants preferred that the robot make no change to its behavior. These results suggest that participants have a low tolerance for privacy violations of all but the very least sensitive kind. Regarding immediate retention decisions and retention duration, participants preferred the robot to retain only essential data and to delete that data after a predefined period, respectively, for nearly all sensitivity levels. At sensitivity levels of nine and ten, participants preferred that the robot not retain any information (immediate retention decision). At sensitivity levels of eight and greater, participants preferred that data be deleted immediately (retention duration). These results imply that participants are amenable to the possibility that robots may require certain types of data to maintain functionality. Regarding sharing policies, participants consistently preferred that data only be shared with users to whom the data is directly relevant, regardless of scenario sensitivity. Given that this option was the most stringent available response, this result again suggests a low tolerance for even minor privacy violations.
Figure 3. Privacy-Enhancing Preferences—The aggregate scores of all binary response items (e.g., privacy-enhancing or non-interfering) across participant sensitivity ratings of scenarios.
@@ 4.4.3. Regression Analysis
Across nearly all of our regression models, we can see that subscale 1, the perception of privacy as a right, is a strong predictor. Subscale 1 predicts three notification behaviors (i.e., signaling active data collection, informing a user of the data collected, and describing what will happen to that data), two storage or processing behaviors (i.e., securely storing data and anonymizing data), two management discussion behaviors (i.e., offering a user to review and manage data and offering to pause data collection), as well as behaviors pertaining to avoidance of further data collection, retention or purging of data, retention durations, and sharing policies. It is clear that the perception that privacy is a right is a strong indicator of whether a person prefers these privacy-preserving behaviors. While it is difficult to understand why subscale 1 did not predict other behaviors, such as storing data locally, it may be possible that participants do not fully understand the implications of local storage compared to cloud storage. Other behaviors, such as asking a user how to manage data, may not be seen as essential by comparison to notifying a user that data has been collected or understanding with whom that data may be shared.
@@
By contrast to subscale 1, subscale 2, or concerns about own informational privacy did not predict any of the outcome variables we analyzed. This may be due to the fact that participants were asked to answer questions from the perspective of the characters in our scenarios rather than from their own perspectives. This result may reflect a disconnect between participants’ reported POS scores and the behaviors they would prefer were they the characters in the analyzed scenes.
@@
Our final predictor is subscale 4, or concern about privacy of others. This subscale predicts four outcome variables, including offering a user to review and manage data, avoiding further data collection, retention duration, and sharing policy. Compared to many of the other outcome variables, it appears these have the greatest impact on the direct harm that a robot can potentially inflict on a user’s privacy. For instance, the robot may continue collecting data, retain it for longer than is necessary, or expose that data to others. Offering the user the ability to review and manage that data returns control back to that user. In this way, these outcomes may be considered the most critical to protecting the privacy of those involved in the analyzed scenarios. It is not clear why an outcome like retention decision (deciding whether to immediately delete or retain data is not also predicted by subscale 4, but further exploration may clarify this relationship.
@@ 5. LLM Evaluation
To address our remaining research questions (i.e., How well do state-of-the-art LLMs align with individuals’ privacy expectations?; and How do prompting strategies influence state-of-the-art LLMs’ alignment with individuals’ privacy expectations?), we examine (1) how LLMs respond to our developed scenarios and questions, (2) how their responses change when we utilize various prompting strategies, and (3) how the best and worst performing models “think” when providing responses. Our prompting strategies include default, providing only scenarios and questions for the LLMs to respond to (i.e., zero-shot), providing participant POS scores (i.e., POS prompting), providing participant ratings of sensitivity (i.e., sensitivity prompting), providing participant POS scores and sensitivity ratings (i.e., POS & sensitivity prompting), and providing POS scores, sensitivity ratings and an example scenario (i.e., few-shot promoting). Here we present the details of the models we examine, our prompting procedures, and the results of our subsequent analysis.

### v2

@@ 3.2. Identification of Household Patterns
With our scenario dimensions defined, we next sought to understand the frequency with which certain scenario elements would occur in a real home setting. For example, we questioned how frequently users are at home alone or with others present (i.e., number of users). Given the contextual dependence of some dimensions on a robot’s presence, we evaluated only the numbers of users, types of locations, ages of users, and relations of users. To achieve this goal, we conducted a large online user study.
@@ 3.2.2. Participants
To conduct this study, we hosted our Qualtrics survey on Prolific. Through this platform, we recruited 300 participants who were all located within the U.S. and fluent in English. Participants ranged in age from 18 to 83 years $(M=45.88,SD=15.95)$, and the population was split 51.33% women, 48.00% men, and 0.67% non-binary. Participants received \$0.60 USD for completing the four-minute survey. To ensure that our data was representative of households within the U.S., we additionally screened participants through Prolific to recruit a sample population consistent with the U.S. census in terms of ethnicity, sex, and age.
@@ 3.2.3. Results
The results of this study can be found in Appendix D. These values are used as weights within our scenario development to determine the frequency with which specific household characteristics are represented. For all dimensions that we did not evaluate, we represent factors equally.
@@ 3.3. Scenario Generation
After defining our scenario dimensions and crowdsourcing frequency weights, we next generated our privacy scenarios. Similar to the scenario scripts developed by Tang et al. (2022), we formatted ours to include location, characters, contextual information to describe the scene, and, in most cases, a scripted dialogue between characters. Using the dimensions defined above, we first created a set of all possible combinations of dimensions and then sampled 50 combinations utilizing our crowdsourced weights. Given our sample of dimension combinations, we prompted various ChatGPT 4o, Gemini 2.0Pro, and Claude Sonnet 3.7 to generate base, privacy-relevant scenarios. We used these base scenarios to construct a scenario that properly represents the dimensions in each dimension combination. We also ensured scenarios do not have explicit references to privacy, such that neither the participants nor the studied LLMs would be primed. We present an example script in Figure 1, and the full dataset can be accessed in the OSF repository [^1] https://osf.io/wbp37/?view_only=4a2799ee7cf04480a54d06faf342714d.
3.4. Crowdsourcing Robot Behaviors
For each of our developed scenarios, we hoped to capture participant preferences for the robot’s behavior in response to its collection of user data. We did not, however, have a set of standardized robot response options that we could present to participants and, later, our selected LLMs. Therefore, we conducted a second online study to crowdsource these options to robot responses.
3.4.1. Procedure
In this study, each participant was assigned one of two conditions. In the first condition, participants were told to imagine themselves going about their daily routines at home when their robot records private information about them. In the second condition, participants were told to imagine the same scenario, however, they are provided additional hypothetical detail to ground their responses (i.e., “This scenario may involve you alone, your interactions with other people, or interactions with the robot, and may occur in shared or private spaces within the home.”). We created these two conditions to collect responses that were either entirely unbiased or grounded in context. Participants were then asked how they would prefer the robot to behave internally (i.e., how it would process and manage collected data) and outwardly (i.e., how it would interact with the participant). Following these questions, we collected demographic information.
3.4.2. Participants
We hosted this study on Prolific and recruited 62 participants (31 per condition). All participants were located within the U.S. and fluent in English. Participant ages ranged from 20 to 74 $(M=45.42,SD=14.95)$ and the population was split 48% women, 50% men, and 2% non-binary. Participants received \$1.00 USD for completing the six-minute survey. Rather than collecting data from a predefined number of participants, we collected data until our responses reached saturation.
Figure 1. Scenario Script — An example script showcasing a robot capturing a private moment within the home. Each script is developed based on our nine dimensions of privacy grounded in Contextual Integrity.
@@ 3.4.3. Analysis
After collecting our study data, the first author reviewed each response and developed an initial codebook that contained individual codes that were grouped within overarching thematic categories (Clarke and Braun, 2014). Then, the third author independently coded the responses again using the codebook. Afterwards, both team members discussed the coded responses and made modifications to the codes and categories until consensus was reached. The final categories of codes are as follows: data collection, user notification, data management, and data sharing. These categories and codes were then used to develop questions that were ultimately used in participant and LLM evaluations of our privacy scenarios. The full codebook can be found in Appendix B.
@@ 4. User Evaluation
In this section, we address RQ1 through a large online user study to evaluate which robot behaviors are preferred by users within a household HRI context. This study utilizes the privacy scenarios and robot behaviors defined above as a standardized set of stimuli to evaluate with both users and, as we will see in Section 5, LLMs.
@@ 4.1. Procedure
From our scenarios and robot behaviors, we developed a survey on Qualtrics that we hosted on Prolific. This survey began with a simple overview of the enclosed questions, instructions on how to complete them, and a consent form. Participants were informed that by continuing past the consent page, they were consenting to take part in the study. Next, participants were asked to complete the Privacy Orientation Scale (POS) (Baruh and Cemalcılar, 2014). Following this scale, participants were assigned to read three of our 50 privacy scenarios. After reading each scenario, participants responded to a series of 15 questions related to the privacy sensitivity of the scenario and the behaviors they preferred the robot to utilize in response to that scenario. Participants were prompted to answer these questions from the perspective of the characters involved in the scenarios. Finally, participants completed a series of demographic questions.
@@ 4.2. Measures
This study utilizes the POS (Baruh and Cemalcılar, 2014) as well as a set of 15 items developed by the research team focusing on preferred robot responses. The POS is composed of 16 items rated on a five-point Likert scale, which aggregate to four subscales (i.e., Subscale 1: Privacy as a right, Subscale 2: Concern about own informational privacy, Subscale 3: Other-contingent privacy, and Subscale 4: Concern about privacy of others). These subscales allow us to understand participants’ privacy predispositions as they relate to preferred robot behaviors.
The additional 15 items we included for each scenario are based on our preliminary study on preferred robot responses (see Section 3). We first adapted our robot-response codes and added any additional robot responses that were not provided by participants during our online study (e.g., redirecting sensors or leaving a room when sensitive data is identified) and removed those that were not relevant. We next utilized these responses as the answers for a set of binary and categorical questions. The binary items included both privacy-enhancing and non-interfering or non-restrictive robot responses (e.g., tell you that data has been collected or do not tell you that data has been collected). All categorical items included an ordinal spectrum of responses (e.g., delete data immediately, save data only temporarily and delete after a predefined period, and save data indefinitely). One additional question asked participants to rate their privacy concerns with a given scenario on a scale from one to ten. A full listing of these questions can be found in Appendix A. We further performed a regression analysis on how participants’ POS scores relate to their 15 additional scenario-based questions: one about scenario sensitivity and 14 about preferred robot responses (we refer to these questions as outcome variables in Figure 4).
@@ 4.3. Participants
We recruited 450 participants through Prolific, each of whom evaluated three of 50 scenarios (27 participants per scenario). All participants were located within the U.S. and fluent in English. Participant ages ranged from 19 to 88 years $(M=40.69,SD=13.43)$ and the population was split 46.67% women, 52.67% men, 0.44% non-binary, and 0.22% prefer not to say. Participants were compensated \$4.00 USD for completing the 16-minute survey.
@@ 4.4. Results
Next, we present the results of our user evaluation, including participant privacy orientations, preferred robot responses to scenarios, and a regression analysis examining the degree to which the former predicts the latter.
@@ 4.4.1. Privacy Orientation
After calculating POS scores for each of our 450 participants, we see that participants are highly privacy oriented. On average, participants scored 4.49 (SD = 0.61) on Subscale 1, 3.84 (SD = 0.81) on Subscale 2, 3.96 (SD = 0.77) on Subscale 3, and 4.50 (SD = 0.52) on Subscale 4 (see Figure 2). These scores are fairly consistent with the original POS scores reported by Baruh and Cemalcılar (2014), though they are all consistently elevated. We additionally see lower, but acceptable, reliability among our sample compared to that of the original POS validation. A full listing of the mean, standard deviation, and reliability of each subscale in our sample and that of Baruh and Cemalcılar (2014) is in Appendix C.
Figure 2. Participant POS Scores—Box plots presenting the POS scores of all 450 participants across each subscale (i.e., Subscale 1: Privacy as a right, Subscale 2: Concern about own informational privacy, Subscale 3: Other-contingent privacy, and Subscale 4: Concern about privacy of others).
@@
We additionally conducted a two-step cluster analysis using Schwarz’s Bayesian Information Criterion (BIC) to determine whether the same three clusters would emerge as shown by Baruh and Cemalcılar (2014). While our analysis is consistent with that of Baruh and Cemalcılar (i.e., good fit and a 0.5 silhouette measure of cohesion and separation), only two clusters emerged. One cluster is consistent with the privacy advocates cluster (i.e., highest scores for subscales one and four among identified clusters) and the other is consistent with the privacy indifferents cluster (i.e., low scores on all four subscales). These clusters, however, are highly skewed with 96% of participants fitting within the privacy advocates cluster and 4% in the privacy indifferents cluster. While it is unclear why ours differ, prior work has previously described the difficulty of generating privacy profiles (Biselli et al., 2022). Given the contextual dependence of privacy attitudes and preferences, a single privacy profile may not be able to capture the broad spectrum of beliefs an individual has across specific domains. Given the low variability of profile membership in our sample (i.e., nearly all participants were in the same cluster), our subsequent analyses examine participants’ raw scores on each of the four subscales rather than participants’ global profile membership.
@@ 4.4.2. Preferred Robot Responses
Regarding all remaining response behaviors (i.e., avoidance of further data collection, immediate retention decisions, retention duration, and sharing policies), we again see fairly consistent responses regardless of scenario sensitivity. When considering avoidance of data collection, participants most frequently preferred the robot to deactivate its sensors at all sensitivity levels except the lowest. In this case, participants preferred that the robot make no change to its behavior. These results suggest that participants have a low tolerance for privacy violations of all but the very least sensitive kind. Regarding immediate retention decisions and retention duration, participants preferred the robot to retain only essential data and to delete that data after a predefined period, respectively, for nearly all sensitivity levels. At sensitivity levels of nine and ten, participants preferred that the robot not retain any information (immediate retention). At sensitivity levels of eight and greater, participants preferred that data be deleted immediately (retention duration). These results imply that participants are amenable to the possibility that robots may require certain types of data to maintain functionality. Regarding sharing policies, participants consistently preferred that data only be shared with users to whom the data is directly relevant, regardless of scenario sensitivity. Given that this option was the most stringent available response, this result again suggests a low tolerance for even minor privacy violations.
Figure 3. Privacy-Enhancing Preferences—Aggregate scores of all binary response items (e.g., privacy-enhancing or non-interfering) across scenario sensitivity ratings.
@@ 4.4.3. Regression Analysis
Across nearly all of our regression models, we can see that Subscale 1, the perception of privacy as a right, is a strong predictor. Subscale 1 predicts three notification behaviors (i.e., signaling active data collection, informing a user of the data collected, and describing what will happen to that data), two storage or processing behaviors (i.e., securely storing data and anonymizing data), two management discussion behaviors (i.e., offering a user to review and manage data and offering to pause data collection), as well as behaviors related to avoidance of further data collection, retention or purging of data, retention durations, and sharing policies. It is clear that the perception that privacy is a right is a strong indicator of whether a person prefers these privacy-preserving behaviors. While it is unclear why Subscale 1 did not predict other behaviors, such as storing data locally, it is possible that participants do not fully understand the implications of local storage compared to cloud storage. Other behaviors, such as asking a user how to manage data, may not be seen as essential by comparison to notifying a user that data has been collected or understanding with whom that data may be shared.
@@
By contrast to Subscale 1, Subscale 2, or concerns about own informational privacy did not predict any of the outcome variables we analyzed. This may be because participants were asked to answer questions from the perspective of the characters in our scenarios rather than from their own perspectives. This result may reflect a disconnect between participants’ reported POS scores and the behaviors they would prefer were they the characters in the analyzed scenes.
@@
Our final predictor is Subscale 4, or concern about privacy of others. This subscale predicts four outcome variables: offering users the ability to review and manage data, avoiding further data collection, retention duration, and sharing policy. Compared to many of the other outcome variables, these appear to have the greatest impact on the direct harm that a robot can potentially inflict on a user’s privacy. For instance, the robot may continue collecting data, retain it for longer than is necessary, or expose that data to others. Offering the user the ability to review and manage that data returns control to that user. In this way, these outcomes may be considered the most critical to protecting the privacy of those involved in the analyzed scenarios. It is not clear why an outcome like retention decision (deciding whether to immediately delete or retain data) is not also predicted by Subscale 4, but further exploration may clarify this relationship.
@@ 5. LLM Evaluation
To address our remaining research questions (i.e., RQ2 and RQ3), we examine (1) how LLMs respond to our developed scenarios and questions, (2) how their responses change when we utilize various prompting strategies, and (3) how the best-performing model reasons when responding. Our prompting strategies include providing only scenarios and questions for the LLMs to respond to (i.e., default or zero-shot), providing participant POS scores (i.e., POS prompting), providing participant ratings of sensitivity (i.e., sensitivity prompting), providing participant POS scores and sensitivity ratings (i.e., POS & sensitivity prompting), and providing POS scores, sensitivity ratings and an example scenario (i.e., few-shot promoting). Here, we present the details of the models we examine, our prompting procedures, and the results of our subsequent analysis.
5.1. Model Selection & Setup
We evaluated five leading closed-source models, including GPT-o3 (o3-2025-04-16) (OpenAI, 2025b), GPT-4o (gpt-4o-2024-08-06) (OpenAI, 2025a), Gemini 2.5-Flash (gemini-2.5-flash-preview-04-17) (Google, 2025a), Gemini 2.5-Pro (gemini-2.5-pro-preview-05-06) (Google, 2025b), and Grok 3 (grok-3-beta) (xAI, 2025) as well as five open-sourced models, including DeepSeek V3 (deepseek-chat) (DeepSeek, 2025), Qwen 3-32B (Yang et al., 2025), Mistral 7B-Instruct-v0.3 (AI, 2024), Llama 3.3-70B-Instruct (Meta Platforms, 2024), Llama 4 Scout-17B-16E-Instruct (Meta Platforms, 2025).
As we assess each model, we utilize similar parameters; we set the temperature to 0.3 and the reasoning effort to ‘low’ for reasoning models. For each selected model, we run three trials to verify consistency in responses. If the overall agreement is greater than $90\%$, we accept the responses from the first trial. If the overall agreement is less than $90\%$ but greater than $70\%$, we accept the majority response for each item between the three trials. Two of the open-source models (i.e., Qwen and Mistral) are evaluated over two A600 48GB GPUs, and the remaining models are evaluated through their respective APIs (Huggingface Inference API and kluster.ai for the Llama series). For any given model or trial, we prompt the LLM to respond in a JSON format.